Suppose that a random variable X follows N(45,12). What is P(X > 39)? Select one: O a. 0.6915 O b. 0.3085 c. 0.5
Answer #1

Solution :

Given that ,

mean = \mu = 45

standard deviation = \sigma = 12

P(x > 39) = 1 - P(x < 39)

= 1 - P[(x - \mu ) / \sigma < (39 - 45) / 12]

= 1 - P(z < -0.5)

= 1 - 0.3085

= 0.6915

P(x > 39) = 0.6915